### v1

# RobotFingerPrint: Unified Gripper Coordinate Space for Multi-Gripper Grasp Synthesis

Abstract
We introduce a novel representation named as the unified gripper coordinate space for grasp synthesis of multiple grippers. The space is a 2D surface of a sphere in 3D using longitude and latitude as its coordinates, and it is shared for all robotic grippers. We propose a new algorithm to map the palm surface of a gripper into the unified gripper coordinate space, and design a conditional variational autoencoder to predict the unified gripper coordinates given an input object. The predicted unified gripper coordinates establish correspondences between the gripper and the object, which can be used in an optimization problem to solve the grasp pose and the finger joints for grasp synthesis. We demonstrate that using the unified gripper coordinate space improves the success rate and diversity in the grasp synthesis of multiple grippers.  [^1] Project page is available at: https://irvlutd.github.io/RobotFingerPrint
I INTRODUCTION
Grasp planning or grasp synthesis is an important problem in robotics. Traditional grasp planners such as GraspIt! [1] and OpenRAVE grasp planner [2] adopt a sampling and evaluation strategy, in which the grasps are sampled around an object and the force closure analysis [3] is utilized to evaluate the grasps sampled. These grasp planners are usually computationally expensive and may require users to specify contact points on the grippers. Recently, learning-based grasp synthesis has received more attention. These methods use generative machine learning models such as variational autoencoders [4, 5] or diffusion models [6] for grasp synthesis. These models are trained on large-scale datasets of grasps with various objects [7, 5, 8]. Therefore, they can predict grasps for unseen objects. However, most learning-based models are only trained for a specific gripper such as the Franka Panda gripper [4, 9] or a five-finger gripper [10]. Consequently, these models cannot support multiple grippers or be used for new grippers outside the training domain.
In order to train a single grasp synthesis model to support multiple grippers, researchers have explored predicting some intermediate representations of grasps, instead of directly predicting grasp pose and joint values which are specific for each gripper. In the literature, two types of intermediate representations of grasps have been explored: gripper keypoints and contact maps. For example, UniGrasp [11] and GeoMatch [12] train neural networks to predict the locations of gripper keypoints on an input object and then solve inverse kinematics to obtain the grasps. GenDexGrasp [13] predicts the contact map of a given object and then solves an optimization problem to generate grasps that comply with the contact map. These intermediate representations can be used with different grippers, even for unseen grippers outside the training domain.
Figure 1: Comparison between three different representations for grasp synthesis: (a) contact map; (b) sparse gripper keypoints; (c) our unified gripper coordinate space
We notice that both gripper keypoints and contact maps have limitations in grasp synthesis. First, gripper keypoints are sparse, i.e., usually a few keypoints are sampled on the gripper. Therefore, noises at keypoint locations can significantly affect grasp quality. In GeoMatch [12], users have to manually specify these keypoints. Second, contact maps are ambiguous for grasp synthesis and their order has to pre-determined. They provide only information about the distances between the grippers and the objects, but there is no information about how each finger should touch the object.
In this work, motivated by the normalized object coordinate space in object pose estimation [14] and the dense human pose estimation [15] from the computer vision community, we introduce a novel intermediate representation of grasps named the Unified Gripper Coordinate Space (UGCS). This space is the two-dimensional surface of a sphere in the three-dimensional space denoted $S^{2}$. The coordinates in this space are longitude $\lambda\in[-\pi,\pi)$ and latitude $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$. This space is shared among any robotic gripper. Given a gripper, we propose an algorithm to map the inner surface of the gripper, i.e., “palm” of the gripper, into UGCS. Therefore, every 3D point on the palm is assigned a 2D coordinate $(\lambda,\varphi)$. We consider these points as the “finger print” of the robotic gripper. In order to synthesize grasps, we train a conditional variational autoencoder (CVAE) to map a latent vector and a point cloud of an object to the unified gripper coordinate space. Specifically, CVAE predicts a unified gripper coordinate $(\lambda,\varphi)$ for every point on the surface of the object. Using the predicted unified gripper coordinates, a correspondence is established between the object surface and the gripper palm. Then we solve an optimization problem using the predicted unified gripper coordinates to obtain the grasp pose and the finger joint values of a grasp.
Compared to using gripper keypoints or contact maps for grasp synthesis, our proposed UGCS establishes dense correspondences between a gripper and an object, which is robust to noises in the CVAE predictions and resolves the ambiguities from contact maps. In addition, since the UGCS is shared among all grippers, it enables grasp transfer between grippers. We conducted grasp synthesis experiments using the MultiDex dataset [5] . Our method, RobotFingerPrint, outperforms GenDexGrasp [5] and GeoMatch [12] in terms of grasp success rate and grasp diversity.
II Related Work
II-A Analytic Grasp Synthesis
Classic grasp planners such as GraspIt! [1] and OpenRAVE grasp planner [2] use analytic approaches to access grasp quality. Such methods employ task wrench space analysis [16, 17] or force closure analysis [3, 18] to assess the feasibility of a proposed grasp. These traditional grasp synthesis methods usually use sampling-based optimization algorithms, such as simulated annealing to sample grasps, and then evaluate them, which makes the optimization quite slow. Some methods also require the user to manually specify the preferred contact regions on the gripper surface prior to starting the grasp generation process. Recent analytic grasp synthesis leverages Differentiable Force Closure (DFC) [19] in grasp optimization, where the optimization runtime is significantly improved. The main limitation of analytic methods is that the success rate of grasping may not be guaranteed. Because force closure analysis cannot account for other factors in grasping such as friction, materials of objects, noises in control, etc.
II-B Learning-based Grasp Synthesis
To improve speed and success in grasp synthesis, more efforts have recently been devoted to learning-based grasp synthesis. Datasets of grasps are first constructed for learning. These grasps are then verified as successful in a physics simulator. Representative datasets include Acronym [7], UniGrasp  [11], DexGraspNet [20], Fast-Grasp’D [8], GenDexGrasp [13] and MultiGripperGrasp [21]. These datasets contain different numbers of grippers, objects, and grasps, which can be used to learn generative models for grasp synthesis such as variational auto-encoders or diffusion models.
Grasp synthesis for a single gripper. Most learning-based methods can only plan grasps for a specific type of robotic gripper. For example, [4, 22, 9] deal with parallel jaw grippers while FFHNet [10] is designed for a five-finger gripper. Unlike analytic approaches, these learning-based methods cannot plan grasps for a variety of robotic grippers.
Grasp synthesis for multiple grippers. Recently, a few approaches are proposed for grasp synthesis of multiple grippers. AdaGrasp [23] uses gripper features to convolve with scene features to generate grasps. UniGrasp [11] trains a neural network to predict finger-tip contact points on objects and then solves inverse kinematics to generate grasps. However, both AdaGrasp and UniGrasp cannot scale-up to grippers with more than 3 fingers. GenDexGrasp [13] trains a conditional variational autoencoder to predict contact map of a given object and then solves an optimization problem using the predicted contact map to generate a grasp. GeoMatch [12] follows a similar idea as UniGrasp [11] to predict keypoints of grippers on object surfaces and then solve inverse kinematics to compute grasps. Different from existing methods, we introduce a novel contact map representation that encodes dense correspondences of gripper surface and object surface. This representation improves success rate and diversity in multi-gripper grasp synthesis across several scenarios.
III Method
III-A Unified Gripper Coordinate Space (UGCS)
The Unified Gripper Coordinate Space (UGCS) is a coordinate space that represents the points on the interior surface of a gripper by mapping them onto the surface of a sphere. This coordinate space is common to different grippers and provides a dense representation of the interior surface. A sphere was chosen because of its smooth surface point representation and the fact that similar contact regions can be marked unambiguously across grippers with varying scales. The UGCS representation assigns a normalized coordinate $(\lambda,\varphi)$, with each component in $[0,1]$ obtained after normalizing the longitude and latitude in their ranges $\lambda\in[-\pi,\pi)$ and $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$. Next, we present the method for constructing such a unified coordinate space across different grippers via their grasps on an associated maximal sphere.
Figure 2: Max spheres across all the 12 different grippers including the gripper set from MultiGripperGrasp [21] dataset.
Figure 3: Maximal Graspable Sphere: We test for multiple spheres with varying radii in parallel via Isaac Sim [24]
III-A1 Maximal Sphere Computation
Since the geometry and kinematics amongst grippers differ significantly, there is no single sphere that can be used to assign this coordinate mapping. Hence, we first created a test bed in simulation to obtain the maximal graspable sphere for each gripper, and used it to assign the coordinate mapping. We used Nvidia’s Isaac Sim [24] robotic simulator based on the MultiGripperGrasp [21] toolkit. The simulation consisted of top-down grasps of spheres with increasing radii. These spheres were initialized with their pole touching the gripper palm, and the position, joint values of the gripper at the moment of contact were recorded. Each grasp is simulated for 3 seconds after gripper comes in contact with the object. A grasp is considered successful if, at the end of the grasp the sphere is still in contact with the gripper. Fig. 3 shows the process of sphere-grasping tests in Isaac Sim. The maximal sphere is then determined by the maximum radius within the subset of successful grasps. The maximal graspable sphere of all grippers in the MultiGripperGrasp [21] and MultiDex [5] datasets can be seen in Fig. 2.
III-A2 UGCS for Gripper Points
Figure 4: Visualization of Unified Gripper Coordinate System on grippers from MultiDex [5] and their corresponding sphere points. The colors indicate where each gripper’s region is mapped on the sphere surface.
We use the maximal sphere for each gripper to assign a $(\lambda,\varphi)$ coordinate to the gripper’s interior surface points. We only consider the interior surface points as they are in contact or close proximity with object for any grasp. This also reduces the ambiguity in which gripper regions are supposed to contact the object, leading to stable grasps. The interior gripper surface points set $P_{G}$ is computed by uniformly sampling a set $R$ of unit ray directions from the sphere center and checking for ray-mesh intersections $\mathcal{I}_{1}(\hat{r},G)$ with the gripper’s meshes.
$P_{G}=\{v_{g}\in G\;|\;\exists\;\hat{r}\in R,\;\text{s.t}\;\mathcal{I}_{1}(\hat{r},G)=v_{g}\}.$
Each intersecting unit ray direction $\hat{r}$ can alternatively be viewed as surface points on a unit sphere with some spherical coordinate $(radius=1;\lambda,\varphi)$. If the unit ray $\hat{r}_{g}$ was used to obtain an interior surface point $v_{g}\in P_{G}$, then we simply assign the ray’s angular component of the spherical coordinate $(\lambda_{\hat{r}_{g}},\varphi_{\hat{r}_{g}})$ as the gripper point’s coordinate: $(\lambda_{v_{g}},\varphi_{v_{g}}):=(\lambda_{\hat{r}_{g}},\varphi_{\hat{r}_{g}})$. Consequently, with a uniform set of rays, we get dense gripper surface points and their associated coordinates. For $M$ gripper interior surface points, we obtain a UGCS gripper representation with shape $(M,2)$ and label this as $\Phi_{G}$. Note that unlike manually selected sparse gripper key points, the UGCS representation is not affected by the ordering of points or the number $M$. Fig. 4 illustrates the idea behind obtaining $\Phi_{G}$ for each gripper.
III-B Grasp Representation
Given a set of grasps on an object with a given gripper, we show how to utilize the UGCS representation with a coordinate map $\Phi$ on the object point cloud. Note that the gripper point cloud $P_{G}$ with its $M$ interior surface points and the corresponding UCGS coordinate map $\Phi_{G}$ are fixed across the different grasps on the object set. Consider an object point cloud $O$ having $N$ points and a grasp for gripper $G$ with its fixed interior surface point set $P_{G}$ in the object frame. Now, we have $N$ object points and $M$ gripper points in the object frame, with each gripper point having a corresponding $(\lambda,\varphi)$ coordinate. For the object’s coordinate map $\Phi_{O}$, we assign coordinates from the closest gripper point.
$\Phi_{O}^{(v_{o})}=\Phi_{G}^{(v_{g}^{*})}\;\text{where}\;v_{g}^{*}=\operatorname*{\arg\!\min}_{v_{g}\in P_{G}}||v_{o}-v_{g}||.$
In this way, we obtain a UGCS coordinate for each object point in $O$. The object map $\Phi_{O}$ provides an alternate grasp representation on the object, and its object-centric form is not constrained by the differences in grippers. A drawback in the above formulation is that it assigns a coordinate $(\lambda,\varphi)$ to each object point $v_{o}$ regardless of its distance to the gripper and without considering whether it is actually in contact or not. In order to represent such no-contact object points in $P_{O}$, we simply map them to a default coordinate of $(\lambda=0,\varphi=0)$.
Object points with such a coordinate give signal to the learning model that they are not in contact with the object. We can safely assign this specific coordinate as we are guaranteed that no gripper point ever gets assigned this coordinate. This specific coordinate represents the opposite pole of the sphere from where it was grasped. The maximal sphere simulation test ensures that no gripper point gets projected to this coordinate (see Fig. 4). Finally, similar to a contact-map dataset on multi-gripper grasps, we create a coordinate-map dataset with grasps on the objects. Fig. 5 shows some examples of such grasp coordinate maps on object surface with the no-contact regions being transparent.
Figure 5: Ground-truth coordinate map $\Phi_{O}$ on the object set with some sample grasps from the MultiDex [5] dataset.
III-C Model Architecture
Figure 6: Method Overview: A randomly sampled latent vector is decoded to a probable object coordinate map $\hat{\Phi}_{O}$. Given a target (possibly unseen) gripper, correspondence is established between object and gripper points using $\hat{\Phi}_{O}$. The same $\hat{\Phi}_{O}$ also guides the pose initialization for the grasp optimization which results in successful grasps.
The overall goal for our RFP method is to synthesize grasps on objects and generalize to unseen grippers and objects. In our modeling setup, we assume that we are given an object point cloud $P_{O}$ with $N$ points. The prediction problem consists of predicting the intermediate grasp representation of object coordinate map $\Phi_{O}$ with a values $(\lambda,\varphi)$ for each of the $N$ points. We pose this as a generative modeling problem and utilize the CVAE architecture used in previous works on hand-object grasping [25, 5].
The goal for the learning model is to reason about the grasps on the object $O$ and predict the intermediate grasp representations in the form of object coordinate maps $\Phi_{O}$. Here each grasp $q$ on $O$ is represented by its associated coordinate map $\Phi_{O}$. Given $P_{O}$ and its coordinate map $\Phi_{O}$, the encoder $p_{\alpha}$ of the C-VAE model learns to encode the object point cloud and the coordinate map into a latent vector $z$ by modeling the distribution $p_{\alpha}(z|P_{O},\Phi_{O})$. The decoder $p_{\beta}$ learns to predict an accurate $\hat{\Phi}_{O}$ on the object given a randomly sampled $z$ by modeling the distribution $p_{\beta}(\Phi_{O}|z,P_{O})$. We utilize PointNet++ [26] based backbone, where the final output consists of $(\hat{\lambda},\hat{\varphi})$ coordinate prediction for each point in the object point cloud.
The model is learned by using a standard loss formulation for CVAE with two components for the training objective function. The first component consists of the reconstruction loss $L_{recon}$ on the ground truth coordinate map $\Phi$ and its prediction $\hat{\Phi}$: $L_{recon}=\sum_{i=1}^{N}||\Phi_{O}^{(i)}-\hat{\Phi}_{O}^{(i)}||$. The second component consists of a KL-divergence term $L_{KL}$ between a prior distribution on the latent space $p_{z}=\mathcal{N}(0,1)$ and the approximate posterior modeled by the decoder $p_{\beta}(\Phi_{O}|z,P_{O})$: $L_{KL}=D_{KL}(p_{z},p_{\beta})$. The total loss is a weighted sum between the two: $L_{total}=w_{1}\;L_{recon}\;+\;w_{2}\;L_{KL}$. During inference, given an object point cloud, we sample a latent vector $z\sim\mathcal{N}(0,1)$ and pass it through the decoder along with the point cloud to obtain the predicted coordinate map $\hat{\Phi}_{O}$. This output is then used as an intermediate representation for grasp optimization on a given gripper as shown in Fig. 6.
III-D Grasp Optimization
III-D1 Coordinate Map Correspondence
The ultimate goal of the method is to produce a grasp configuration $q_{G}$ with pose and joint values on an object $O$ and a target gripper $G$. We generate $q_{G}$ for a predicted object coordinate map $\hat{\Phi}_{O}$ and $G$ via a correspondence-driven optimization with the gripper coordinate map $\Phi_{G}$ on its in interior surface point set $P_{G}$. In contact map based grasp synthesis, the predicted contact map is oblivious to the gripper fingers and hence ambiguous to interpret. A key advantage of UGCS is that we can easily establish a correspondence between an object $O$ with $\hat{\Phi}_{O}$ and gripper $G$ with $\Phi_{G}$. Given $N$ object points, and $M$ gripper points, we establish a correspondence between them by assigning each object point $v_{o}$ to a gripper point $v_{g}^{*(o)}$ based on the great-circle distance $D_{H}$ (also known as the Haversine distance) between their respective coordinates $(\hat{\lambda}_{v_{o}},\hat{\varphi}_{v_{o}})$ and $(\lambda_{v_{g}},\varphi_{v_{g}})$. The distance is computed after projecting the coordinates back from the normalized space to the original ranges for latitude and longitude.
$v_{g}^{*(o)}=\operatorname*{\arg\!\min}_{v_{g}\in P_{G}}\;\;D_{H}\left((\hat{\lambda}_{v_{o}},\hat{\varphi}_{v_{o}})\;;\;(\lambda_{v_{g}},\varphi_{v_{g}})\right).$
Using this procedure we obtain a dense mapping between the object and gripper points and it can guide the grasp optimization towards suitable grasps for $\hat{\Phi}_{O}$. While establishing this correspondence, we do not consider probable no-contact object points which are inferred on the basis of low values for both of their predicted coordinates. The final result from the correspondence are two matching subsets $P_{O_{c}}\subset O$ and $P_{G_{c}}\subset P_{G}$ from the object and gripper points, respectively.
III-D2 Pose Initialization
A possible way to initialize the optimization procedure for $q_{G}$ would be to sample random gripper poses around the object via some heuristics. However, this can lead to suboptimal grasps due to improper initialization. We observe that in addition to the correspondence, the predicted coordinate map $\Phi_{O}$ contains another signal to the optimization procedure in terms of probable gripper pose. During the construction for UGCS, each of the grippers grasps the sphere with their palm touching the sphere’s south pole with a normalized coordinate value $(\lambda=0,\varphi=1)$. As a consequence, on the predicted coordinate map $\hat{\Phi}$, we infer a set $S_{B}$ of probable object points that map to gripper base/palm by thresholding on $\hat{\lambda}$ and $\hat{\varphi}$ using pre-defined thresholds $\lambda_{ub}$ and $\varphi_{lb}$, respectively:
$S_{B}=\{v_{o}\in P_{O}\;s.t\;\hat{\lambda}_{v_{o}}\leq\lambda_{ub}\;\;\text{and}\;\;\hat{\varphi}_{v_{o}}\geq\varphi_{lb}\}.$
The initial gripper position is set to $0.1m$ away from the mean of palm-matching set $S_{B}$ along the mean surface normal for it. The initial gripper 6D pose is determined by aligning the gripper palm normal to the opposite of the mean surface normal from $S_{B}$ as shown in Fig. 6.
TABLE I: Comparison of RFP with prior works on success rate and diversity metrics
Method Success Rate (%) ↑ Diversity (rad) ↑
ezgripper barrett shadowhand Mean ezgripper barrett shadowhand Mean
DFC [19] 58.81 85.48 72.86 72.38 0.3095 0.3770 0.3472 0.3446
AdaGrasp [23] 60.00 80.00 - 70.00 0.0003 0.0002 - 0.0003
GenDexGrasp [5] 43.44 71.72 77.03 64.01 0.2380 0.2480 0.2110 0.2323
GeoMatch [12] 75.00 90.00 72.50 79.17 0.1880 0.2490 0.2050 0.2140
RFP (Ours) 72.34 90.00 76.87 79.73 0.2330 0.4240 0.2520 0.3030
RFP* (w/o refinement) 52.5 86.56 55.47 64.84 0.2326 0.4256 0.2519 0.3034
III-D3 Optimization
The grasp configuration $q_{G}$ is obtained by minimizing an objective function that considers hand-object penetration, joint value validity and hand-object distance. A differentiable kinematics model is used to provide gradients on $q$ using the objective function. Note that the hand-object distance is only computed with the corresponding point pairs from $P_{O_{c}}$ and $P_{G_{c}}$. The optimization problem is defined as
$q_{G}^{*}=\operatorname*{\arg\!\min}_{q}\;E_{dist}(P_{O_{c}},P_{G_{c}};q)+E_{pen}(O,G;q)+E_{n}(q).$
The terms for penetration $E_{pen}$ and joint value validity $E_{n}$ are adopted from [5]. They constrain the grasp $q$ to not collide with the object, and keep the joint values within specified limits. The correspondence-driven $E_{dist}$ acts as a denser version of Inverse Kinematics (IK) by not being limited to only a pre-specified set of sparse keypoints.
IV Experiments
The grasps generated by our method were evaluated using the same simulation protocol used by [5, 12]. Each grasp is tested for 1 second with 60 simulation steps by applying a consistent acceleration of $0.5ms^{-2}$ on the object in the NVIDIA Isaac Gym simulation. This process is repeated six times across the positive and negatives axes of the world coordinate frame. A grasp is considered as a failure if the object moves by 2cm at the end of any test. A grasp is labeled as successful if it passes all six tests. We report the same evaluation metrics of grasp success rate and the diversity in successful grasps (standard deviation of the joint values) in order to be consistent with the comparison. The test set comprised of the same split of 10 unseen objects as in [5, 12], and we sampled 64 grasps for each gripper-object pair for evaluation.
IV-A Grasp Synthesis with Seen Grippers
First, we evaluate the RFP by training with all 5 grippers and evaluating on the test object set. This experiment setting is same as in [12]. We report metrics for Ezgripper, Barrett and Shadowhand as these three grippers have their test results from [5]. Table I shows the success rate and diversity of all the methods using models trained on all the grippers. Results for DFC [19], AdaGrasp [23], GenDexGrasp [5] and GeoMatch [12] are taken from GeoMatch [12]. We observe that RFP outperforms comparable methods in terms of grasp diversity and success rate. Our results are comparable to [12] and better than [5]. It is important to note that while the RFP demonstrates competitive performance with [12], it is capable of dealing with unseen grippers and can generalize well without explicit retraining. Fig. 7 shows some successful grasps from our method.
TABLE II: Performance on unseen grippers for RobotFingerPrint (RFP) and GenDexGrasp (GDG) [5]. RFP* represents the experiment with no-refinement ablation.
Gripper Method Succ. Rate (%) ↑ Diversity (rad) ↑
Ezgripper RFP 63.28 0.224
RFP* 42.97 0.214
GDG 38.59 0.248
Barrett RFP 83.90 0.400
RFP* 77.96 0.396
GDG 70.31 0.267
Shadow RFP 67.96 0.238
RFP* 51.09 0.238
GDG 77.19 0.207
Figure 7: Successful grasps on unseen objects with RFP.
IV-B Grasp Synthesis with Unseen Grippers
Furthermore, we compared the performance of our method with that of GenDexGrasp [5] when using out-of-domain grippers and the same test set of unseen objects. The model for an out-of-domain gripper was trained with grasps from all grippers except itself. The comparison is presented in Table II where we see a strong improvement over the prior work across most of the metrics. RFP under-performs GenDexGrasp with ShadowHand in success rate but shows better diversity. It could be due to difficulty in setting up the correspondence for such a dexterous gripper in presence of noisy coordinate predictions. On the other hand, RFP outperforms in the case of Ezgripper and Barrett as the correspondence helps the coordinate mapping. With GenDexGrasp, the multi-fingered contact maps from other grippers could induce uncertainty for the finger mapping with Ezgripper or Barrett.
Figure 8: Similar grasp prediction across different grippers using the same predicted object coordinate map $\hat{\Phi}_{O}$.
IV-C Ablation Study: Grasp Refinement in Simulation
We carried out an ablation study for the grasp refinement step used in the simulation evaluation tests as in [5, 12]. A contact-aware refinement is applied to the predicted grasp $q$ via a single gradient update step to obtain force closure around the object. By removing this refinement, we wanted to test the efficacy of our method in terms of its original output. We compare it with the refined outputs from other methods under both in-domain and out-domain settings for the grippers. As seen in Tables I and Table II, RFP without refinement remains competitive even under the removal of this refinement step and establishes strong baselines. It even outperforms on some metrics from prior works [5, 19, 23] which include this refinement during their evaluation. The dense correspondence-driven grasp optimization and a stable pose initialization are able to help the optimization, avoiding bad minima and converging to good quality grasps.
Figure 9: Failures from the simulation test due to insufficient force closure or not utilizing all the fingers for grasping.
IV-D Ablation Study: Same Prediction for Different Grippers
Another possible reason for the competitive performance with unseen grippers is visualized in Fig. 8. On the same predicted coordinate map on the object, we observe that the three different grippers grasp the object in similar fashion, especially in terms of approach and position around the object. The mean difference in grasp positions was $0.03m$ for Barrett and Ezgripper while it was around $0.06m$ for both of them with the Shadowhand grasps. This indicates the efficacy of RFP in generalizing the grasps from different grippers in a unified coordinate space. We also present some examples of failure cases in Fig. 9.
V Conclusion and Future Work
We introduce RobotFingerPrint (RFP), a novel and effective way to unify grasp representations across different grippers into a common coordinate space. The unified grasp representation is dense in nature, requires no manual specification and has an unambiguous mapping to each gripper’s individual fingers. Such a mapping helps in synthesizing diverse and stable grasping configurations via a hand-object correspondence and well-initialized grasp optimization which results in strong performance across several evaluation scenarios. We plan to extend the core idea of RFP-driven grasp synthesis to partially observed and incomplete point clouds in real-world grasping scenarios and also utilizing it for grasp transfer from humans. Another avenue of exploration would be to investigate a learned reverse mapping from a coordinate map to grasp prediction without the intermediate grasp optimization.
ACKNOWLEDGMENT
This work was supported by the Sony Research Award Program and the National Science Foundation (NSF) under Grant No. 2346528.

### v2

@@
-RobotFingerPrint: Unified Gripper Coordinate Space for Multi-Gripper Grasp Synthesis
+RobotFingerPrint: Unified Gripper Coordinate Space for Multi-Gripper Grasp Synthesis and Transfer
 Abstract
-We introduce a novel representation named as the unified gripper coordinate space for grasp synthesis of multiple grippers.
-The space is a 2D surface of a sphere in 3D using longitude and latitude as its coordinates, and it is shared for all robotic grippers.
-We propose a new algorithm to map the palm surface of a gripper into the unified gripper coordinate space, and design a conditional variational autoencoder to predict the unified gripper coordinates given an input object.
-The predicted unified gripper coordinates establish correspondences between the gripper and the object, which can be used in an optimization problem to solve the grasp pose and the finger joints for grasp synthesis.
-We demonstrate that using the unified gripper coordinate space improves the success rate and diversity in the grasp synthesis of multiple grippers.
-[^1] Project page is available at: https://irvlutd.github.io/RobotFingerPrint
+We introduce a novel grasp representation named the Unified Gripper Coordinate Space (UGCS) for grasp synthesis and grasp transfer.
+Our representation leverages spherical coordinates to create a shared coordinate space across different robot grippers, enabling it to synthesize and transfer grasps for both novel objects and previously unseen grippers.
+The strength of this representation lies in the ability to map palm and fingers of a gripper and the unified coordinate space.
+Grasp synthesis is formulated as predicting the unified spherical coordinates on object surface points via a conditional variational autoencoder.
+The predicted unified gripper coordinates establish exact correspondences between the gripper and object points, which is used to optimize grasp pose and joint values.
+Grasp transfer is facilitated through the point-to-point correspondence between any two (potentially unseen) grippers and solved via a similar optimization.
+Extensive simulation and real-world experiments showcase the efficacy of the unified grasp representation for grasp synthesis in generating stable and diverse grasps.
+Similarly, we showcase real-world grasp transfer from human demonstrations across different objects.
+[^1] Project page with code and videos is available at: https://irvlutd.github.io/RobotFingerPrint
 I INTRODUCTION
-Grasp planning or grasp synthesis is an important problem in robotics.
+Grasp synthesis is a fundamental problem in robot manipulation.
 Traditional grasp planners such as GraspIt!
-[1] and OpenRAVE grasp planner [2] adopt a sampling and evaluation strategy, in which the grasps are sampled around an object and the force closure analysis [3] is utilized to evaluate the grasps sampled.
-These grasp planners are usually computationally expensive and may require users to specify contact points on the grippers.
-Recently, learning-based grasp synthesis has received more attention.
-These methods use generative machine learning models such as variational autoencoders [4, 5] or diffusion models [6] for grasp synthesis.
-These models are trained on large-scale datasets of grasps with various objects [7, 5, 8].
-Therefore, they can predict grasps for unseen objects.
-However, most learning-based models are only trained for a specific gripper such as the Franka Panda gripper [4, 9] or a five-finger gripper [10].
-Consequently, these models cannot support multiple grippers or be used for new grippers outside the training domain.
-In order to train a single grasp synthesis model to support multiple grippers, researchers have explored predicting some intermediate representations of grasps, instead of directly predicting grasp pose and joint values which are specific for each gripper.
-In the literature, two types of intermediate representations of grasps have been explored: gripper keypoints and contact maps.
+[1] and OpenRAVE [2] adopt a sampling and evaluation strategy, in which the grasps are sampled around an object and a force closure analysis [3] is used to evaluate the grasps.
+These grasp planners are usually computationally expensive and may require users to manually specify contact points on the grippers.
+Recently, learning-based grasp synthesis has received more attention for generalization to objects and grippers.
+These methods use generative machine learning models such as variational autoencoders [4, 5] or diffusion models [6] for grasp generation.
+They are trained on large-scale grasp datasets with various objects, which enable them to predict grasps for unseen objects [7, 5, 8].
+However, most learning-based models are trained only for a specific gripper [4, 9, 10].
+Consequently, these models lack the ability to synthesize grasps for robotic hands outside their training domain.
+In order to train a single grasp synthesis model to support multiple grippers, researchers have explored predicting intermediate grasp representations instead of directly predicting the grasp pose and joint values which are specific for each gripper.
+In the literature, two types of intermediate grasp representation have been developed: gripper keypoints and contact maps.
 For example, UniGrasp [11] and GeoMatch [12] train neural networks to predict the locations of gripper keypoints on an input object and then solve inverse kinematics to obtain the grasps.
-GenDexGrasp [13] predicts the contact map of a given object and then solves an optimization problem to generate grasps that comply with the contact map.
-These intermediate representations can be used with different grippers, even for unseen grippers outside the training domain.
+Furthermore, GenDexGrasp [5] predicts a contact map of a given object and then solves an optimization problem to generate grasps that comply with the generated contact map.
+These intermediate grasps representations have been used to synthesize grasps for various grippers, even for unseen grippers outside the training domain.
 Figure 1: Comparison between three different representations for grasp synthesis: (a) contact map; (b) sparse gripper keypoints; (c) our unified gripper coordinate space
-We notice that both gripper keypoints and contact maps have limitations in grasp synthesis.
-First, gripper keypoints are sparse, i.e., usually a few keypoints are sampled on the gripper.
-Therefore, noises at keypoint locations can significantly affect grasp quality.
-In GeoMatch [12], users have to manually specify these keypoints.
-Second, contact maps are ambiguous for grasp synthesis and their order has to pre-determined.
-They provide only information about the distances between the grippers and the objects, but there is no information about how each finger should touch the object.
-In this work, motivated by the normalized object coordinate space in object pose estimation [14] and the dense human pose estimation [15] from the computer vision community, we introduce a novel intermediate representation of grasps named the Unified Gripper Coordinate Space (UGCS).
-This space is the two-dimensional surface of a sphere in the three-dimensional space denoted $S^{2}$.
-The coordinates in this space are longitude $\lambda\in[-\pi,\pi)$ and latitude $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$.
-This space is shared among any robotic gripper.
-Given a gripper, we propose an algorithm to map the inner surface of the gripper, i.e., “palm” of the gripper, into UGCS.
-Therefore, every 3D point on the palm is assigned a 2D coordinate $(\lambda,\varphi)$.
-We consider these points as the “finger print” of the robotic gripper.
-In order to synthesize grasps, we train a conditional variational autoencoder (CVAE) to map a latent vector and a point cloud of an object to the unified gripper coordinate space.
-Specifically, CVAE predicts a unified gripper coordinate $(\lambda,\varphi)$ for every point on the surface of the object.
-Using the predicted unified gripper coordinates, a correspondence is established between the object surface and the gripper palm.
-Then we solve an optimization problem using the predicted unified gripper coordinates to obtain the grasp pose and the finger joint values of a grasp.
-Compared to using gripper keypoints or contact maps for grasp synthesis, our proposed UGCS establishes dense correspondences between a gripper and an object, which is robust to noises in the CVAE predictions and resolves the ambiguities from contact maps.
-In addition, since the UGCS is shared among all grippers, it enables grasp transfer between grippers.
-We conducted grasp synthesis experiments using the MultiDex dataset [5] .
-Our method, RobotFingerPrint, outperforms GenDexGrasp [5] and GeoMatch [12] in terms of grasp success rate and grasp diversity.
+We illustrate the limitations in both gripper keypoints and contact maps for grasp synthesis in Fig.
+1.
+On one hand, gripper keypoints are sparse.
+Therefore, noise for any predicted keypoint location can significantly affect grasp quality.
+Users must manually specify these keypoints [12], and the keypoint prediction order is predetermined, thus inhibiting flexibility.
+Similarly, contact maps are ambiguous for grasp synthesis when mapping grippers to contact locations.
+They only provide information about the distances between the gripper and the object, but there is no information on how and where the gripper fingers should touch the object [5].
+Consequently, these representations are neither suitable for transferring one grasp from one gripper to another gripper.
+Motivated by the normalized object coordinate space in object pose estimation [13] and dense human pose estimation [14], we introduce a novel intermediate grasp representation named the Unified Gripper Coordinate Space (UGCS).
+This space, denoted as $S^{2}$, is the two-dimensional surface of a sphere in a three-dimensional space.
+Any point within the UGCS is defined by its longitude $\lambda\in[-\pi,\pi)$ and its latitude $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$.
+This space is intended to be shared between all robotic hands and used to map the inner surface of the grippers (gripper “palm”).
+Here, every 3D point on the palm is assigned a 2D coordinate $(\lambda,\varphi)$.
+We refer to these points as the “finger print” of the robotic gripper.
+For grasp synthesis, we train a Conditional Variational Auto-Encoder (CVAE) to map a latent vector and a point cloud of an object to the UGCS.
+Specifically, CVAE predicts the UGCS points $(\lambda,\varphi)$ for every point on the surface of the object.
+Then, an objective function is optimized using the predicted coordinates to obtain the synthesized grasp configuration.
+Cross-gripper grasp transfer is computed by first initializing a mutual correspondence between surface points for source and target grippers, via their respective coordinate maps.
+Then the target gripper grasp is inferred from an optimization that tries to minimize the distance between the corresponding points.
+Compared to keypoints and contact maps as grasp representations, our proposed UGCS establishes dense correspondences between a gripper and an object.
+We conducted grasp synthesis experiments using the MultiDex dataset [5] and compared it to methods that use alternative intermediate grasp representations [5, 12].
+Our method, RobotFingerPrint (RFP), was found to compare favorably against prior works in terms of both grasp success rate and grasp diversity.
+Furthermore, we tested our method with out-of-domain Fetch robot [15] gripper by performing real-world manipulation on the SceneReplica [16] benchmark.
+The grasps generated by the RFP were able to outperform on the model-based grasp scenario with the GraspIt [1] baseline of the benchmark.
+We also showed the effectiveness in grasp transfer from human hand to the Fetch gripper.
+Such a transfer pipeline can be an effective tool in learning from demonstration scenarios with purely visual information.
+Our contributions are summarized as follows.
+  • We introduce the unified gripper coordinate space that establishes dense correspondences across different grippers via a novel shared representation space.
+  • Diverse and stable object-centric grasp synthesis across in/out-domain grippers and object sets with stronger pose initialization.
+  • Generalized grasp transfer across different grippers without requiring manual re-targeting.
 II Related Work
 II-A Analytic Grasp Synthesis
-Classic grasp planners such as GraspIt!
+Classical grasp planners such as GraspIt!
 [1] and OpenRAVE grasp planner [2] use analytic approaches to access grasp quality.
@@
 Recent analytic grasp synthesis leverages Differentiable Force Closure (DFC) [20] in grasp optimization, where the optimization runtime is significantly improved.
-The main limitation of analytic methods is that the success rate of grasping may not be guaranteed.
-Because force closure analysis cannot account for other factors in grasping such as friction, materials of objects, noises in control, etc.
+The main limitation of analytic methods is that the success rate of grasping may not be guaranteed because force closure analysis cannot account for other factors in grasping such as friction, materials of objects and noises in control.
 II-B Learning-based Grasp Synthesis
-To improve speed and success in grasp synthesis, more efforts have recently been devoted to learning-based grasp synthesis.
-Datasets of grasps are first constructed for learning.
-These grasps are then verified as successful in a physics simulator.
+To improve speed and success in grasp synthesis, recent efforts have been devoted to learning-based grasp synthesis.
+Datasets have been designed for machine learning methods on grasping via grasp stability verification in physics simulators.
 Representative datasets include Acronym [7], UniGrasp [11], DexGraspNet [21], Fast-Grasp’D [8], GenDexGrasp [5] and MultiGripperGrasp [22].
 These datasets contain different numbers of grippers, objects, and grasps, which can be used to learn generative models for grasp synthesis such as variational auto-encoders or diffusion models.
-Grasp synthesis for a single gripper.
+II-B1 Single-Gripper Models
 Most learning-based methods can only plan grasps for a specific type of robotic gripper.
 For example, [4, 23, 9] deal with parallel jaw grippers while FFHNet [10] is designed for a five-finger gripper.
-Unlike analytic approaches, these learning-based methods cannot plan grasps for a variety of robotic grippers.
-Grasp synthesis for multiple grippers.
-Recently, a few approaches are proposed for grasp synthesis of multiple grippers.
-AdaGrasp [23] uses gripper features to convolve with scene features to generate grasps.
-UniGrasp [11] trains a neural network to predict finger-tip contact points on objects and then solves inverse kinematics to generate grasps.
+Unlike classical approaches, these learning-based methods cannot plan grasps for a variety of robotic hands.
+II-B2 Multi-Gripper Models
+In recent years, a few approaches have been proposed for grasp synthesis of multiple grippers.
+For example, AdaGrasp [24] convolves over both normalized gripper features and scene features to predict grasps, which in turn enables it to generalize over robotic hands.
+UniGrasp [11] trains a neural network to predict finger-tip contact points on objects, and then solves inverse kinematics to generate grasps.
 However, both AdaGrasp and UniGrasp cannot scale-up to grippers with more than 3 fingers.
-GenDexGrasp [13] trains a conditional variational autoencoder to predict contact map of a given object and then solves an optimization problem using the predicted contact map to generate a grasp.
+On the other hand, GenDexGrasp [5] use a CVAE to predict contact maps on a given object and then solve an optimization problem using the predicted contact map to generate a grasp.
 GeoMatch [12] follows a similar idea as UniGrasp [11] to predict keypoints of grippers on object surfaces and then solve inverse kinematics to compute grasps.
-Different from existing methods, we introduce a novel contact map representation that encodes dense correspondences of gripper surface and object surface.
-This representation improves success rate and diversity in multi-gripper grasp synthesis across several scenarios.
-III Method
-III-A Unified Gripper Coordinate Space (UGCS)
-The Unified Gripper Coordinate Space (UGCS) is a coordinate space that represents the points on the interior surface of a gripper by mapping them onto the surface of a sphere.
+We introduce a new method for multi-gripper grasp synthesis based on our UGCS representation.
+II-C Grasp Transfer
+For grasp transfer, previous work either focuses on the problem of cross-object transfer [25, 26] or formulates it as retrieval from an offline dataset [27] that inhibits generalization to unseen objects and grippers.
+Unlike existing methods, we introduce a novel grasp representation that encodes dense correspondences of gripper surfaces and object surfaces.
+This representation enables grasp transfer that generalizes to novel grippers and objects.
+Figure 2: Max spheres across all the 12 different grippers including the gripper set from the MultiGripperGrasp [22] dataset.
+III Unified Gripper Coordinate Space
+The Unified Gripper Coordinate Space (UGCS) is a coordinate space that maps the points on the interior surface of a gripper onto the surface of a sphere.
 This coordinate space is common to different grippers and provides a dense representation of the interior surface.
@@
 The UGCS representation assigns a normalized coordinate $(\lambda,\varphi)$, with each component in $[0,1]$ obtained after normalizing the longitude and latitude in their ranges $\lambda\in[-\pi,\pi)$ and $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$.
-Next, we present the method for constructing such a unified coordinate space across different grippers via their grasps on an associated maximal sphere.
-Figure 2: Max spheres across all the 12 different grippers including the gripper set from MultiGripperGrasp [21] dataset.
-Figure 3: Maximal Graspable Sphere: We test for multiple spheres with varying radii in parallel via Isaac Sim [24]
-III-A1 Maximal Sphere Computation
-Since the geometry and kinematics amongst grippers differ significantly, there is no single sphere that can be used to assign this coordinate mapping.
-Hence, we first created a test bed in simulation to obtain the maximal graspable sphere for each gripper, and used it to assign the coordinate mapping.
-We used Nvidia’s Isaac Sim [24] robotic simulator based on the MultiGripperGrasp [21] toolkit.
-The simulation consisted of top-down grasps of spheres with increasing radii.
-These spheres were initialized with their pole touching the gripper palm, and the position, joint values of the gripper at the moment of contact were recorded.
-Each grasp is simulated for 3 seconds after gripper comes in contact with the object.
-A grasp is considered successful if, at the end of the grasp the sphere is still in contact with the gripper.
-Fig.
+Next, we present the method for constructing such a unified coordinate space across any gripper via their grasp on an associated maximal sphere.
+Figure 3: Maximal Graspable Sphere: We test for multiple spheres with varying radii in parallel using Isaac Sim [28]
+III-A Maximal Sphere Computation
+Since the geometry and kinematics between grippers differ significantly, there is no single sphere that can be used to assign this coordinate mapping.
+Hence, we created a test bed in simulation to obtain the maximal graspable sphere of each gripper for assigning the coordinate mapping.
+We used the NVIDIA Isaac Sim [28] robotic simulator where the simulation consisted of top-down grasps of spheres with increasing radii.
+These spheres were initialized with a common pole position and each gripper performed an encompassing grasp from the south pole.
+Each grasp was simulated for 3 seconds after the gripper came in contact with the sphere.
+The maximal spheres of grippers in the MultiGripperGrasp [22] and MultiDex [5] datasets can be seen in Fig.
+2, while Fig.
 3 shows the process of sphere-grasping tests in Isaac Sim.
-The maximal sphere is then determined by the maximum radius within the subset of successful grasps.
-The maximal graspable sphere of all grippers in the MultiGripperGrasp [21] and MultiDex [5] datasets can be seen in Fig.
-2.
-III-A2 UGCS for Gripper Points
+III-B UGCS for Gripper Points
 Figure 4: Visualization of Unified Gripper Coordinate System on grippers from MultiDex [5] and their corresponding sphere points.
 The colors indicate where each gripper’s region is mapped on the sphere surface.
-We use the maximal sphere for each gripper to assign a $(\lambda,\varphi)$ coordinate to the gripper’s interior surface points.
-We only consider the interior surface points as they are in contact or close proximity with object for any grasp.
-This also reduces the ambiguity in which gripper regions are supposed to contact the object, leading to stable grasps.
-The interior gripper surface points set $P_{G}$ is computed by uniformly sampling a set $R$ of unit ray directions from the sphere center and checking for ray-mesh intersections $\mathcal{I}_{1}(\hat{r},G)$ with the gripper’s meshes.
-$P_{G}=\{v_{g}\in G\;|\;\exists\;\hat{r}\in R,\;\text{s.t}\;\mathcal{I}_{1}(\hat{r},G)=v_{g}\}.$
-Each intersecting unit ray direction $\hat{r}$ can alternatively be viewed as surface points on a unit sphere with some spherical coordinate $(radius=1;\lambda,\varphi)$.
-If the unit ray $\hat{r}_{g}$ was used to obtain an interior surface point $v_{g}\in P_{G}$, then we simply assign the ray’s angular component of the spherical coordinate $(\lambda_{\hat{r}_{g}},\varphi_{\hat{r}_{g}})$ as the gripper point’s coordinate: $(\lambda_{v_{g}},\varphi_{v_{g}}):=(\lambda_{\hat{r}_{g}},\varphi_{\hat{r}_{g}})$.
-Consequently, with a uniform set of rays, we get dense gripper surface points and their associated coordinates.
-For $M$ gripper interior surface points, we obtain a UGCS gripper representation with shape $(M,2)$ and label this as $\Phi_{G}$.
+We use the maximal sphere for each gripper to assign a $(\lambda,\varphi)$ coordinate to the interior surface points of the gripper.
+Interior surface points are considered as they are close to the objects.
+This also reduces the ambiguity in which gripper points are supposed to contact the object, leading to stable grasps.
+The interior gripper surface points set $P_{G}$ is computed by uniformly sampling a set $R$ of unit ray directions from the sphere center and checking for ray-mesh intersection $\mathcal{I}_{1}(\hat{\mathbf{r}},G)$ with the mesh $G$ of the gripper:
+$P_{G}=\{\mathbf{v}_{g}\in G\;|\;\exists\;\hat{\mathbf{r}}\in R,\;\text{s.t.,}\;\mathcal{I}_{1}(\hat{\mathbf{r}},G)=\mathbf{v}_{g}\},$ (1)
+Here $\mathbf{v}_{g}$ is the intersection point of the ray $\hat{\mathbf{r}}$ and the gripper mesh $G$.
+Each intersecting unit ray direction $\hat{\mathbf{r}}$ is also viewed as a surface point on a unit sphere with spherical coordinate $(\lambda,\varphi)$.
+We then assign the ray’s angular component of the spherical coordinate as the gripper point’s coordinate: $(\lambda_{\mathbf{v}_{g}},\varphi_{\mathbf{v}_{g}}):=(\lambda_{\hat{\mathbf{r}}},\varphi_{\hat{\mathbf{r}}})$.
+With a uniform set of rays, we obtain dense gripper surface points and their associated coordinates.
+For $M$ gripper points, we obtain the UGCS representation with shape $(M,2)$ and denote this as $\Phi_{G}$.
 Note that unlike manually selected sparse gripper key points, the UGCS representation is not affected by the ordering of points or the number $M$.
@@
 4 illustrates the idea behind obtaining $\Phi_{G}$ for each gripper.
-III-B Grasp Representation
-Given a set of grasps on an object with a given gripper, we show how to utilize the UGCS representation with a coordinate map $\Phi$ on the object point cloud.
-Note that the gripper point cloud $P_{G}$ with its $M$ interior surface points and the corresponding UCGS coordinate map $\Phi_{G}$ are fixed across the different grasps on the object set.
-Consider an object point cloud $O$ having $N$ points and a grasp for gripper $G$ with its fixed interior surface point set $P_{G}$ in the object frame.
-Now, we have $N$ object points and $M$ gripper points in the object frame, with each gripper point having a corresponding $(\lambda,\varphi)$ coordinate.
-For the object’s coordinate map $\Phi_{O}$, we assign coordinates from the closest gripper point.
-$\Phi_{O}^{(v_{o})}=\Phi_{G}^{(v_{g}^{*})}\;\text{where}\;v_{g}^{*}=\operatorname*{\arg\!\min}_{v_{g}\in P_{G}}||v_{o}-v_{g}||.$
-In this way, we obtain a UGCS coordinate for each object point in $O$.
+The gripper point cloud $P_{G}\;;|P_{G}|=M$ and its UGCS coordinate map $\Phi_{G}$ are fixed for the gripper across its different grasps on objects.
+Figure 5: Ground-truth coordinate map $\Phi_{O}$ on the object set with some sample grasps from the MultiDex [5] dataset.
+III-C Grasp Representation
+In this section, we detail how to represent grasps in an object-centric manner using the UGCS coordinate map $\Phi$ on the object’s point cloud.
+Consider an object $O$ ’s point cloud $P_{O}$ and a grasp from gripper $G$ with its fixed surface point set $P_{G}$ and coordinates $(\lambda,\varphi)$ for each gripper point.
+To construct the coordinate map $\Phi_{O}$ of the object, we assign coordinates from the closest gripper point:
+$\Phi_{O}^{(\mathbf{v}_{o})}=\Phi_{G}^{(\mathbf{v}_{g}^{*})},\;\text{where}\;\mathbf{v}_{g}^{*}=\operatorname*{\arg\!\min}_{\mathbf{v}_{g}\in P_{G}}||\mathbf{v}_{o}-\mathbf{v}_{g}||,$ (2)
+Here $\mathbf{v}_{o}$ indicates an object point.
+In this way, we obtain the UGCS coordinate for each object point.
 The object map $\Phi_{O}$ provides an alternate grasp representation on the object, and its object-centric form is not constrained by the differences in grippers.
-A drawback in the above formulation is that it assigns a coordinate $(\lambda,\varphi)$ to each object point $v_{o}$ regardless of its distance to the gripper and without considering whether it is actually in contact or not.
-In order to represent such no-contact object points in $P_{O}$, we simply map them to a default coordinate of $(\lambda=0,\varphi=0)$.
-Object points with such a coordinate give signal to the learning model that they are not in contact with the object.
-We can safely assign this specific coordinate as we are guaranteed that no gripper point ever gets assigned this coordinate.
-This specific coordinate represents the opposite pole of the sphere from where it was grasped.
-The maximal sphere simulation test ensures that no gripper point gets projected to this coordinate (see Fig.
+For non-contact object points in $P_{O}$ (determined by a threshold of 1cm on gripper-object distance), we simply map them to a default coordinate of $(\lambda=0,\varphi=0)$.
+This is the sphere’s north pole with original coordinates as $(-\pi/2,-\pi)$, and is never mapped to a gripper point as a consequence of maximal sphere test (see Fig.
 4).
@@
 5 shows some examples of such grasp coordinate maps on object surface with the no-contact regions being transparent.
-Figure 5: Ground-truth coordinate map $\Phi_{O}$ on the object set with some sample grasps from the MultiDex [5] dataset.
-III-C Model Architecture
+IV Grasp Synthesis
+IV-A Model Architecture
 Figure 6: Method Overview: A randomly sampled latent vector is decoded to a probable object coordinate map $\hat{\Phi}_{O}$.
 Given a target (possibly unseen) gripper, correspondence is established between object and gripper points using $\hat{\Phi}_{O}$.
-The same $\hat{\Phi}_{O}$ also guides the pose initialization for the grasp optimization which results in successful grasps.
-The overall goal for our RFP method is to synthesize grasps on objects and generalize to unseen grippers and objects.
+$\hat{\Phi}_{O}$ also guides the pose initialization for the grasp optimization which results in successful grasps.
+One of the goals for RFP is to synthesize grasps on in/out of domain objects and grippers.
 In our modeling setup, we assume that we are given an object point cloud $P_{O}$ with $N$ points.
-The prediction problem consists of predicting the intermediate grasp representation of object coordinate map $\Phi_{O}$ with a values $(\lambda,\varphi)$ for each of the $N$ points.
-We pose this as a generative modeling problem and utilize the CVAE architecture used in previous works on hand-object grasping [25, 5].
-The goal for the learning model is to reason about the grasps on the object $O$ and predict the intermediate grasp representations in the form of object coordinate maps $\Phi_{O}$.
-Here each grasp $q$ on $O$ is represented by its associated coordinate map $\Phi_{O}$.
-Given $P_{O}$ and its coordinate map $\Phi_{O}$, the encoder $p_{\alpha}$ of the C-VAE model learns to encode the object point cloud and the coordinate map into a latent vector $z$ by modeling the distribution $p_{\alpha}(z|P_{O},\Phi_{O})$.
-The decoder $p_{\beta}$ learns to predict an accurate $\hat{\Phi}_{O}$ on the object given a randomly sampled $z$ by modeling the distribution $p_{\beta}(\Phi_{O}|z,P_{O})$.
-We utilize PointNet++ [26] based backbone, where the final output consists of $(\hat{\lambda},\hat{\varphi})$ coordinate prediction for each point in the object point cloud.
-The model is learned by using a standard loss formulation for CVAE with two components for the training objective function.
-The first component consists of the reconstruction loss $L_{recon}$ on the ground truth coordinate map $\Phi$ and its prediction $\hat{\Phi}$: $L_{recon}=\sum_{i=1}^{N}||\Phi_{O}^{(i)}-\hat{\Phi}_{O}^{(i)}||$.
-The second component consists of a KL-divergence term $L_{KL}$ between a prior distribution on the latent space $p_{z}=\mathcal{N}(0,1)$ and the approximate posterior modeled by the decoder $p_{\beta}(\Phi_{O}|z,P_{O})$: $L_{KL}=D_{KL}(p_{z},p_{\beta})$.
-The total loss is a weighted sum between the two: $L_{total}=w_{1}\;L_{recon}\;+\;w_{2}\;L_{KL}$.
-During inference, given an object point cloud, we sample a latent vector $z\sim\mathcal{N}(0,1)$ and pass it through the decoder along with the point cloud to obtain the predicted coordinate map $\hat{\Phi}_{O}$.
+The prediction problem consists of inferring an intermediate grasp representation of the object coordinate map $\Phi_{O}$ with values $(\lambda,\varphi)$ for $N$ points.
+We pose this as a generative modeling problem and utilize the CVAE architecture used in previous work on hand-object grasping [29, 5].
+The goal of the learning model is to reason about the grasps on the object $O$ and predict the grasp representations in the form of object coordinate maps $\Phi_{O}$ for each grasp on it.
+Given $P_{O}$ and its coordinate map $\Phi_{O}$, the encoder $p_{\alpha}$ of the CVAE model learns to encode the object point cloud and the coordinate map into a latent vector $\mathbf{z}$ by modeling the distribution $p_{\alpha}(\mathbf{z}|P_{O},\Phi_{O})$.
+The decoder $p_{\beta}$ learns to predict an accurate $\hat{\Phi}_{O}$ on the object given a randomly sampled $\mathbf{z}$ by modeling the distribution $p_{\beta}(\Phi_{O}|\mathbf{z},P_{O})$.
+We utilize a PointNet++ [30] based backbone, where the final output consists of $(\hat{\lambda},\hat{\varphi})$ coordinate prediction for each point in the object point cloud.
+The model is trained using a standard loss formulation for CVAE with two components for the objective function.
+The first component is the reconstruction loss $L_{\text{recon}}$ on the ground truth coordinate map $\Phi$ and its prediction $\hat{\Phi}$: $L_{\text{recon}}=\sum_{i=1}^{N}||\Phi_{O}^{(i)}-\hat{\Phi}_{O}^{(i)}||$.
+The second component consists of a KL-divergence term $L_{\text{KL}}$ between a prior distribution on the latent space $p_{z}=\mathcal{N}(0,\mathcal{I})$ and the approximate posterior modeled by the decoder $p_{\beta}(\Phi_{O}|\mathbf{z},P_{O})$: $L_{\text{KL}}=D_{\text{KL}}(p_{z},p_{\beta})$.
+The total loss is a weighted sum between the two: $L_{\text{total}}=\;L_{\text{recon}}\;+\;w_{\text{KL}}\;L_{\text{KL}}$, where the weight $w_{\text{KL}}$ is annealed over training epochs.
+During inference, given an object point cloud, we sample a latent vector $\mathbf{z}\sim\mathcal{N}(0,\mathcal{I})$ and pass it through the decoder with the object point cloud to obtain the predicted coordinate map $\hat{\Phi}_{O}$.
 This output is then used as an intermediate representation for grasp optimization on a given gripper as shown in Fig.
 6.
-III-D Grasp Optimization
-III-D1 Coordinate Map Correspondence
-The ultimate goal of the method is to produce a grasp configuration $q_{G}$ with pose and joint values on an object $O$ and a target gripper $G$.
-We generate $q_{G}$ for a predicted object coordinate map $\hat{\Phi}_{O}$ and $G$ via a correspondence-driven optimization with the gripper coordinate map $\Phi_{G}$ on its in interior surface point set $P_{G}$.
-In contact map based grasp synthesis, the predicted contact map is oblivious to the gripper fingers and hence ambiguous to interpret.
-A key advantage of UGCS is that we can easily establish a correspondence between an object $O$ with $\hat{\Phi}_{O}$ and gripper $G$ with $\Phi_{G}$.
-Given $N$ object points, and $M$ gripper points, we establish a correspondence between them by assigning each object point $v_{o}$ to a gripper point $v_{g}^{*(o)}$ based on the great-circle distance $D_{H}$ (also known as the Haversine distance) between their respective coordinates $(\hat{\lambda}_{v_{o}},\hat{\varphi}_{v_{o}})$ and $(\lambda_{v_{g}},\varphi_{v_{g}})$.
-The distance is computed after projecting the coordinates back from the normalized space to the original ranges for latitude and longitude.
-$v_{g}^{*(o)}=\operatorname*{\arg\!\min}_{v_{g}\in P_{G}}\;\;D_{H}\left((\hat{\lambda}_{v_{o}},\hat{\varphi}_{v_{o}})\;;\;(\lambda_{v_{g}},\varphi_{v_{g}})\right).$
-Using this procedure we obtain a dense mapping between the object and gripper points and it can guide the grasp optimization towards suitable grasps for $\hat{\Phi}_{O}$.
-While establishing this correspondence, we do not consider probable no-contact object points which are inferred on the basis of low values for both of their predicted coordinates.
-The final result from the correspondence are two matching subsets $P_{O_{c}}\subset O$ and $P_{G_{c}}\subset P_{G}$ from the object and gripper points, respectively.
-III-D2 Pose Initialization
-A possible way to initialize the optimization procedure for $q_{G}$ would be to sample random gripper poses around the object via some heuristics.
-However, this can lead to suboptimal grasps due to improper initialization.
-We observe that in addition to the correspondence, the predicted coordinate map $\Phi_{O}$ contains another signal to the optimization procedure in terms of probable gripper pose.
-During the construction for UGCS, each of the grippers grasps the sphere with their palm touching the sphere’s south pole with a normalized coordinate value $(\lambda=0,\varphi=1)$.
-As a consequence, on the predicted coordinate map $\hat{\Phi}$, we infer a set $S_{B}$ of probable object points that map to gripper base/palm by thresholding on $\hat{\lambda}$ and $\hat{\varphi}$ using pre-defined thresholds $\lambda_{ub}$ and $\varphi_{lb}$, respectively:
-$S_{B}=\{v_{o}\in P_{O}\;s.t\;\hat{\lambda}_{v_{o}}\leq\lambda_{ub}\;\;\text{and}\;\;\hat{\varphi}_{v_{o}}\geq\varphi_{lb}\}.$
-The initial gripper position is set to $0.1m$ away from the mean of palm-matching set $S_{B}$ along the mean surface normal for it.
-The initial gripper 6D pose is determined by aligning the gripper palm normal to the opposite of the mean surface normal from $S_{B}$ as shown in Fig.
+IV-B Grasp Optimization
+IV-B1 Coordinate Map Correspondence
+The final output for grasp synthesis should be a grasp configuration $\mathbf{q}_{G}$, with gripper pose and joint values.
+We generate $\mathbf{q}_{G}$ for a predicted object $O$’s coordinate map $\hat{\Phi}_{O}$ and target gripper $G$ via a UGCS correspondence optimization.
+In contact map based grasp synthesis, predicted contact map is oblivious to the gripper fingers and hence ambiguous to interpret.
+A key advantage of our method is that we can easily establish a correspondence between object and gripper points via the coordinate maps.
+Given $N$ object points, and $M$ gripper points, correspondence is found by assigning each object point $\mathbf{v}_{o}$ to a gripper point $\mathbf{v}_{g}^{*(o)}$.
+Assignment is based on the great-circle distance $D_{H}$ (Haversine distance) between their respective coordinates $(\hat{\lambda}_{\mathbf{v}_{o}},\hat{\varphi}_{\mathbf{v}_{o}})$ and $(\lambda_{\mathbf{v}_{g}},\varphi_{\mathbf{v}_{g}})$:
+$\mathbf{v}_{g}^{*(o)}=\operatorname*{\arg\!\min}_{\mathbf{v}_{g}\in P_{G}}\;\;D_{H}\left((\hat{\lambda}_{\mathbf{v}_{o}},\hat{\varphi}_{\mathbf{v}_{o}})\;;\;(\lambda_{\mathbf{v}_{g}},\varphi_{\mathbf{v}_{g}})\right).$ (3)
+This establishes a dense mapping with two matching subsets of $P_{O_{c}}$ and $P_{G_{c}}$ from the object and gripper points.
+IV-B2 Pose Initialization
+Randomly sampling gripper poses around the object for initialization like in prior works can lead to suboptimal grasps as the optimization gets stuck in local minima.
+In addition to the correspondence, the predicted coordinate map $\hat{\Phi}_{O}$ contains another signal to the optimization in terms of gripper pose initialization.
+By definition, in UGCS, each gripper grasps its sphere with palm/base touching the sphere’s south pole with a normalized coordinate value $(\lambda=0,\varphi=1)$.
+Using this fact, we infer a set $S_{B}$ of probable object points that map to the gripper base/palm region.
+$S_{B}$ contains the points whose predicted coordinates $(\hat{\lambda},\hat{\varphi})\in\hat{\Phi}_{O}$ are within some thresholds $\lambda_{\text{ub}}$ and $\varphi_{\text{lb}}$ for base pole’s coordinates:
+$S_{B}=\{\mathbf{v}_{o}\in P_{O}\;s.t.,\;\hat{\lambda}_{\mathbf{v}_{o}}\leq\lambda_{\text{ub}}\;\;\text{and}\;\;\hat{\varphi}_{\mathbf{v}_{o}}\geq\varphi_{\text{lb}}\}.$ (4)
+The initial gripper position is set to $0.1m$ away from the mean location along the mean surface normal for palm-matching point set $S_{B}$.
+The approach direction for the gripper palm is opposite to the surface normal as shown in Fig.
 6.
@@
 RFP* (w/o refinement) 52.5 86.56 55.47 64.84 0.2326 0.4256 0.2519 0.3034
-III-D3 Optimization
-The grasp configuration $q_{G}$ is obtained by minimizing an objective function that considers hand-object penetration, joint value validity and hand-object distance.
-A differentiable kinematics model is used to provide gradients on $q$ using the objective function.
-Note that the hand-object distance is only computed with the corresponding point pairs from $P_{O_{c}}$ and $P_{G_{c}}$.
-The optimization problem is defined as
-$q_{G}^{*}=\operatorname*{\arg\!\min}_{q}\;E_{dist}(P_{O_{c}},P_{G_{c}};q)+E_{pen}(O,G;q)+E_{n}(q).$
-The terms for penetration $E_{pen}$ and joint value validity $E_{n}$ are adopted from [5].
-They constrain the grasp $q$ to not collide with the object, and keep the joint values within specified limits.
-The correspondence-driven $E_{dist}$ acts as a denser version of Inverse Kinematics (IK) by not being limited to only a pre-specified set of sparse keypoints.
-IV Experiments
-The grasps generated by our method were evaluated using the same simulation protocol used by [5, 12].
-Each grasp is tested for 1 second with 60 simulation steps by applying a consistent acceleration of $0.5ms^{-2}$ on the object in the NVIDIA Isaac Gym simulation.
-This process is repeated six times across the positive and negatives axes of the world coordinate frame.
-A grasp is considered as a failure if the object moves by 2cm at the end of any test.
-A grasp is labeled as successful if it passes all six tests.
-We report the same evaluation metrics of grasp success rate and the diversity in successful grasps (standard deviation of the joint values) in order to be consistent with the comparison.
-The test set comprised of the same split of 10 unseen objects as in [5, 12], and we sampled 64 grasps for each gripper-object pair for evaluation.
-IV-A Grasp Synthesis with Seen Grippers
-First, we evaluate the RFP by training with all 5 grippers and evaluating on the test object set.
-This experiment setting is same as in [12].
-We report metrics for Ezgripper, Barrett and Shadowhand as these three grippers have their test results from [5].
+IV-B3 Optimization
+The grasp configuration $\mathbf{q}_{G}$ is computed by minimizing an objective function that considers hand-object penetration, joint value validity, and hand-object distance.
+A differentiable kinematics gripper model provides gradients on $\mathbf{q}$.
+Here, the hand-object distance $E_{\text{dist}}$ is computed only with corresponding point pairs from $P_{O_{c}}$ and $P_{G_{c}}$.
+$\mathbf{q}_{G}^{*}=\operatorname*{\arg\!\min}_{\mathbf{q}}E_{\text{dist}}(P_{O_{c}},P_{G_{c}};\mathbf{q})+E_{\text{p}}({O},G;\mathbf{q})+E_{\text{n}}(\mathbf{q}).$ (5)
+The terms for hand-object collision $E_{\text{p}}$ and gripper joint validity $E_{\text{n}}$ are adopted from [5].
+They constrain the grasp $\mathbf{q}$ to avoid collisions, and keep the joint values within limits.
+The correspondence-driven $E_{\text{dist}}$ acts as a denser version of inverse kinematics with sparse keypoints seen in prior works.
+Figure 7: Grasp Transfer Pipeline: Human demonstration to executing transferred grasp on Fetch
+V Grasp Transfer
+The unified gripper coordinate space also enables grasp transfer between any two (potentially unseen) grippers without retargeting.
+Any new, unseen gripper is fit to UGCS via its maximal sphere and subsequently obtaining its surface point coordinate map.
+Suppose we are given two grippers, $s$ (source) and $t$ (target) with their coordinate maps $\Phi_{s},\Phi_{t}$.
+In the grasp transfer problem, we are given a grasp $\mathbf{q}_{s}$ for the source and we need to infer a similar grasp $\mathbf{q}_{t}$ for the target.
+$\mathbf{q}_{t}$ is inferred via the grasp transfer optimization shown in Eq.
+(6).
+Here we do not assume access to object point cloud for the source gripper demonstration and consequently, do not involve hand-object point correspondence and collision terms.
+The two grasps $\mathbf{q}_{s},\mathbf{q}_{t}$ give us the transformed gripper point clouds $P_{s},P_{t}$ by applying the grasps to respective grippers.
+A set of mutually corresponding pairs of points $P_{s}^{(c)},P_{t}^{(c)}$ via their respective coordinate maps $\Phi_{t},\Phi_{s}$ is computed using a similar formulation seen in Eq.
+(3).
+Given the corresponding set, we enforce a distance similarity between the point sets so that the target gripper points align with the source gripper.
+$\mathbf{q}_{t}^{*}=\operatorname*{\arg\!\min}_{\mathbf{q}}\;E_{\text{dist}}(P_{s}^{(c)},P_{t}^{(c)};\mathbf{q})+E_{\text{n}}(\mathbf{q}),$ (6)
+where $E_{\text{dist}}$ and $E_{\text{n}}$ are defined as in Eq.
+(5).
+Note that going from dexterous grippers to simple, 2-finger gripper is not well defined for plain contact-map driven transfer.
+The dense correspondences from our method resolve this ambiguity and output a functional and similar grasping pose.
+An illustration of the grasp transfer pipeline between human hand and Fetch robot gripper is shown in Fig.
+7.
+VI Experiments
+VI-A Grasp Synthesis in Simulation
+We used the MultiDex dataset [5] to train and evaluate our models in simulation making it comparable to prior works [5, 12] and their evaluation protocol.
+Each grasp generated by our method is tested for 6 seconds with 60 simulation steps per second and by applying a consistent acceleration of $0.5ms^{-2}$ on the object in the NVIDIA Isaac Gym simulation.
+A grasp is considered as failure if the object moves by 2cm at the end of any test else it is labeled as successful.
+We record the same evaluation metrics of grasp success rate and the diversity in successful grasps (standard deviation of joint values) in order to be consistent with the comparison.
+The test set was comprised of the same split of 10 unseen objects as in [5, 12], and we sampled 64 grasps for each gripper-object pair for evaluation.
+VI-A1 Grasp Synthesis with Seen Grippers
+We evaluated the RFP by training with all 5 grippers and evaluating on the test object set.
+We report metrics for Ezgripper, Barrett and Shadowhand according to [5, 12].
 Table I shows the success rate and diversity of all the methods using models trained on all the grippers.
@@
 Our results are comparable to [12] and better than [5].
-It is important to note that while the RFP demonstrates competitive performance with [12], it is capable of dealing with unseen grippers and can generalize well without explicit retraining.
+It is important to note that while the RFP demonstrates competitive performance with [12], it is capable of dealing with unseen grippers whereas [12] cannot do so.
 Fig.
-7 shows some successful grasps from our method.
+8 shows some successful and failure grasps of our method.
 TABLE II: Performance on unseen grippers for RobotFingerPrint (RFP) and GenDexGrasp (GDG) [5]. RFP* represents the experiment with no-refinement ablation.
@@
 RFP* 42.97 0.214
-GDG 38.59 0.248
+GDG [5] 38.59 0.248
 Barrett RFP 83.90 0.400
 RFP* 77.96 0.396
-GDG 70.31 0.267
+GDG [5] 70.31 0.267
 Shadow RFP 67.96 0.238
 RFP* 51.09 0.238
-GDG 77.19 0.207
-Figure 7: Successful grasps on unseen objects with RFP.
-IV-B Grasp Synthesis with Unseen Grippers
+GDG [5] 77.19 0.207
+Figure 8: Simulation Results: (a) Successful grasps on unseen objects with RFP ; (b) Some failure modes: insufficient force closure or not using all fingers
+VI-A2 Grasp Synthesis with Unseen Grippers
 Furthermore, we compared the performance of our method with that of GenDexGrasp [5] when using out-of-domain grippers and the same test set of unseen objects.
@@
 Figure 9: Similar grasp prediction across different grippers using the same predicted object coordinate map $\hat{\Phi}_{O}$.
-IV-C Ablation Study: Grasp Refinement in Simulation
+Figure 10: Illustration of our real-world grasping pipeline. Our method is used to synthesize grasps given a 3D object model.
+Figure 11: Grasp Transfer Experiments: Examples of huamn demo (from robot’s view) with success/failure cases.
+VI-B Ablation Studies
+VI-B1 Grasp Refinement during Simulation Test
 We carried out an ablation study for the grasp refinement step used in the simulation evaluation tests as in [5, 12].
-A contact-aware refinement is applied to the predicted grasp $q$ via a single gradient update step to obtain force closure around the object.
+A contact-aware refinement is applied to the predicted grasp $\mathbf{q}$ via a single gradient update step to obtain force closure around the object.
 By removing this refinement, we wanted to test the efficacy of our method in terms of its original output.
@@
 The dense correspondence-driven grasp optimization and a stable pose initialization are able to help the optimization, avoiding bad minima and converging to good quality grasps.
-Figure 9: Failures from the simulation test due to insufficient force closure or not utilizing all the fingers for grasping.
-IV-D Ablation Study: Same Prediction for Different Grippers
-Another possible reason for the competitive performance with unseen grippers is visualized in Fig.
-8.
-On the same predicted coordinate map on the object, we observe that the three different grippers grasp the object in similar fashion, especially in terms of approach and position around the object.
-The mean difference in grasp positions was $0.03m$ for Barrett and Ezgripper while it was around $0.06m$ for both of them with the Shadowhand grasps.
+VI-B2 Grasp Inference on Same Object Coordinate Map
+On the same predicted coordinate map of the object, we observe that the three different grippers grasp the object in a similar fashion, especially in terms of approach and position around the object as visualized in Fig.
+9.
+The mean difference in grasp positions was $0.03m$ for Barrett and Ezgripper while it was circa $0.06m$ for both of them with the Shadowhand grasps.
 This indicates the efficacy of RFP in generalizing the grasps from different grippers in a unified coordinate space.
-We also present some examples of failure cases in Fig.
-9.
-V Conclusion and Future Work
+VI-C Real World Grasping Evaluation
+We used the SceneReplica [16] real-world benchmark to evaluate the grasps generated by our method.
+SceneReplica benchmark is comprised of 20 standardized cluttered scenes with 5 YCB objects each.
+It benchmarks the task of grasping each object in scene and moving to a dropoff location with 100 trials in total.
+Grasping and pick-place success rates were the evaluation metrics used for the trials.
+An example setup of the experiments is shown in Fig.
+10.
+For our model-based pipeline, we utilized the GDRNPP model [31] for object pose estimation from the robot’s RGB-D camera.
+Grasp generation was performed for each object using RFP by inferring coordinate maps on the object using the trained model, and optimizing for the Fetch gripper’s grasp configuration.
+The MultiGripperGrasp [22] toolkit was used to rank the grasps based on a fall time evaluation metric in simulation.
+Moveit was used for motion planning to evaluate and execute collision-free grasps in descending order of evaluation scores.
+It is worth noting that of these 16 YCB objects, 6 objects for which we are generating grasps are unseen to our model.
+Moreover, the Fetch robot’s 2 finger gripper is also a new gripper outside of the model’s training domain.
+The grasp was deemed successful if the robot lifted the object, with the pick-and-place task failing only if the object fell before placement.
+As seen in Table III, the grasps computed by RFP outperform those given by GraspIt!
+on the benchmark.
+Despite the training dataset not containing grasps from an equivalent 2-finger gripper, RFP is able to synthesize reasonable real-world grasps for the Fetch gripper.
+Please see supplementary video for grasping examples.
+TABLE III: Grasping and Pick-and-Place (P&P) success rates on the SceneReplica benchmark [16] for 100 trials.
+Perception Module Grasp Module P&P (%) ↑ Grasping (%)↑
+GDRNPP [31] RFP (ours) 70 73
+GDRNPP [31] GraspIt! [1] 66 69
+PoseRBPF [32] GraspIt! [1] 58 64
+VI-D Human-to-Robot Grasp Transfer
+A key advantage of the RFP over previous works is the ability to fit novel and unseen grippers at test time via the unified coordinate map, e.g.
+from a human hand model.
+We evaluated grasp transfer from humans by collecting human grasp demonstrations and executing the transferred grasps with a Fetch robot gripper, shown in Fig.
+7.
+For a human demonstration, we infer the MANO [33] hand parameters with the HaMeR [34] model.
+Using the sphere coordinate mapping defined in Sec.
+III-B, we have the correspondences for human model to the Fetch gripper used for grasp transfer.
+We evaluated the grasp transfer pipeline in real world on 14 YCB objects from SceneReplica [16] by making the Fetch robot grasp and lift the object using the transfer method.
+We achieve a grasping success rate of $0.785$ ($11/14$), with examples shown in Fig.
+11.
+Most failures stem from an incorrect initialization from noisy human hand pose inference.
+VII Conclusion and Future Work
 We introduce RobotFingerPrint (RFP), a novel and effective way to unify grasp representations across different grippers into a common coordinate space.
-The unified grasp representation is dense in nature, requires no manual specification and has an unambiguous mapping to each gripper’s individual fingers.
-Such a mapping helps in synthesizing diverse and stable grasping configurations via a hand-object correspondence and well-initialized grasp optimization which results in strong performance across several evaluation scenarios.
-We plan to extend the core idea of RFP-driven grasp synthesis to partially observed and incomplete point clouds in real-world grasping scenarios and also utilizing it for grasp transfer from humans.
-Another avenue of exploration would be to investigate a learned reverse mapping from a coordinate map to grasp prediction without the intermediate grasp optimization.
-ACKNOWLEDGMENT
-This work was supported by the Sony Research Award Program and the National Science Foundation (NSF) under Grant No. 2346528.
+The unified grasp representation requires no manual specification and has a precise, dense mapping for each gripper’s individual fingers.
+This helps in synthesizing diverse and stable grasp configurations via (1) hand-object correspondence and (2) well-initialized grasp pose.
+RFP shows strong performance across several evaluation scenarios in both simulation and real-world testing with unseen Fetch gripper.
+Further, it enables direct grasp transfer between novel grippers without manual labeling and enables downstream applications in learning from human demonstrations.
+Future work would be to address a potential limitation in collision-aware grasp generation on partial object observations, and real-time grasp inference for use in a closed-loop setting.
+Acknowledgment. This work was supported by the Sony Research Award Program and the National Science Foundation (NSF) under Grant No. 2346528.